# Port-Hamiltonian Neural ODE Networks on Lie Groups For Robot Dynamics Learning and Control

### v1

@@ Abstract
Accurate models of robot dynamics are critical for safe and stable control and generalization to novel operational conditions. Hand-designed models, however, may be insufficiently accurate, even after careful parameter tuning. This motivates the use of machine learning techniques to approximate the robot dynamics over a training set of state-control trajectories. The dynamics of many robots are described in terms of their generalized coordinates on a matrix Lie group, e.g. on $SE(3)$ for ground, aerial, and underwater vehicles, and generalized velocity, and satisfy conservation of energy principles. This paper proposes a (port-)Hamiltonian formulation over a Lie group of the structure of a neural ordinary differential equation (ODE) network to approximate the robot dynamics. In contrast to a black-box ODE network, our formulation guarantees energy conservation principle and Lie group’s constraints by construction and explicitly accounts for energy-dissipation effect such as friction and drag forces in the dynamics model. We develop energy shaping and damping injection control for the learned, potentially under-actuated Hamiltonian dynamics to enable a unified approach for stabilization and trajectory tracking with various robot platforms.
@@ I Introduction
Motion planning and optimal control algorithms depend on the availability of accurate system dynamics models. Models obtained from first principles and calibrated over a small set of parameters via system identification [1] are widely used for unmanned ground vehicles (UGVs), unmanned aerial vehicles (UAVs), and unmanned underwater vehicles (UUVs). Such models may over-simplify or even incorrectly describe the underlying structure of the dynamical system, leading to bias and modeling errors that cannot be corrected by adjusting a few parameters. Data-driven techniques [2, 3, 4, 5, 6] have emerged as a powerful approach to approximate system dynamics with an over-parameterized machine learning model, trained over a dataset of system state and control trajectories. Neural networks are expressive function approximation models, capable of identifying and generalizing interaction patterns from the training data. Training neural network models, however, typically requires large amounts of data and computation time, which may be impractical in mobile robotics applications. Recent works [7, 8, 9, 10, 11, 12, 13, 14] have considered a hybrid approach, where prior knowledge of the physics, governing the system dynamics, is used to assist the learning process. The dynamics of physical systems obey kinematic constraints and energy conservation laws. These laws are known to be universally true but a black-box machine learning model might struggle to infer them from the training data, causing poor generalization. Instead, prior knowledge may be encoded into the learning model, e.g., using a prior distribution [3], a graph-network forward kinematic model [15], or a network architecture reflecting the structure of Lagrangian [16] or Hamiltonian [10] mechanical systems. Moreover, many physical robot platforms are composed of rigid-body interconnections and their state evolution respects the structure of a Lie group [17], e.g., the position and orientation kinematics of a rigid body evolve on the $SE(3)$ Lie group [18].
@@
Lagrangian and Hamiltonian mechanics [19, 20] provide physical system descriptions that can be integrated into the structure of a neural network [10, 21, 11, 22, 23, 24]. Prior work, however, has only considered vector-valued states, when designing Lagrangian- or Hamiltonian-structured neural networks. This limits the applicability of these techniques as many common robot systems have states on a Lie group. For example, Hamiltonian equations of motion are available for orientation but existing formulations rely predominantly on $3$ dimensional vector parametrizations, such as Euler angles [25, 26], which suffer from singularities. In our preliminary work [27], we design a neural ordinary differential equation (ODE) network [28], whose structure captures Hamiltonian dynamics over the $SE(3)$ manifold [29]. Our model guarantees by construction that long-term trajectory predictions satisfy $SE(3)$ constraints and conserve total energy. Inspired by [23, 30], we model kinetic energy and potential energy by separate neural networks, each governed by a set of Hamiltonian equations on $SE(3)$. The Hamiltonian formulation can be generalized to a Port-Hamiltonian one [31], enabling us to design an energy-based controller for trajectory tracking. Our preliminary work [27] is developed specifically for the $SE(3)$ manifold and does not consider dissipation elements that drain energy from the system, such as friction or drag forces. In this paper, we generalize our port-Hamiltonian neural ODE network to embed general matrix Lie group constraints and introduce an energy dissipation term, represented by another neural network, to model friction and air drag in physical systems. We verify our approach extensively with simulated robot systems, including a pendulum and a Crazyflie quadrotor, and with several real quadrotor platforms. In summary, this paper makes the following contributions.
  • We design a neural ODE model that respects port-Hamiltonian dynamics over a matrix Lie group to enable data-driven learning of rigid-body system dynamics.
@@
  • We demonstrate our dynamics learning and control approach with simulated robot systems (a pendulum and a quadrotor) and several real quadrotor robots.
@@ II Related Work
Hamiltonian-based methods [10, 21, 11, 22, 23, 24, 14] use a Hamiltonian formulation [19, 20] of the system dynamics, instead, in terms of generalized coordinates $\mathbf{{\boldsymbol{\mathfrak{q}}}}$, generalized momenta $\mathbf{{\boldsymbol{\mathfrak{p}}}}$, and a Hamiltonian function, $\mathcal{H}(\mathbf{{\boldsymbol{\mathfrak{q}}}},\mathbf{{\boldsymbol{\mathfrak{p}}}})$, representing the total energy of the system. Greydanus et al. [10] model the Hamiltonian as a neural network and update its parameters by minimizing the discrepancy between its symplectic gradients and the time derivatives of the states $(\mathbf{{\boldsymbol{\mathfrak{q}}}},\mathbf{{\boldsymbol{\mathfrak{p}}}})$. This approach, however, requires that the state time derivatives are available in the training data set. Chen et al. [11], Zhong et al. [23] relax this assumption by using differentiable leapfrog integrators [45] and differentiable ODE solvers [28], respectively. The need for time derivatives of the states is eliminated by back-propagating a loss function measuring state discrepancy through the ODE solvers via the adjoint method. Our work extends the approach in [23, 30] by formulating the Hamiltonian dynamics over a matrix Lie group, which enforces kinematic constraints in the neural ODE network used to learn the dynamics. Toth et al. [46] and Mason et al. [47] show that, instead of from state trajectories, the Hamiltonian function can be learned from high-dimensional image observations. Finzi et al. [22] show that using Cartesian coordinates with explicit constraints improves both the accuracy and data efficiency for the Lagrangian- and Hamiltonian-based approaches. In a closely related work, Zhong et al. [30] showed that dissipating elements, such as friction or air drag, can be incorporated in a Hamiltonian-based neural ODE network by reformulating the system dynamics in port-Hamiltonian form [31]. The continuous-time equations of motions in Lagrangian or Hamiltonian dynamics can also be discretized using variational integrators [48] to learn discrete-time Lagrangian and Hamiltonian systems [49, 50, 51, 52] and provide long-term prediction for control methods such as model predictive control [53]. This approach eliminates the need to use an ODE solver to roll out the dynamics but its prediction accuracy depends on the discretization time step.
While most existing dynamics learning approaches focus on Euclidean dynamics, many robot systems have states evolving on a matrix Lie group. We develop a neural ODE framework that incorporates both the law of energy conservation, via a Hamiltonian formulation, and the matrix Lie group constraints in the model. Furthermore, few dynamics learning papers consider control design based on the learned model. We develop a general trajectory tracking controller for Lie group Hamiltonian dynamics.
@@
The Hamiltonian formulation and its port-Hamiltonian generalization [31] are built around the notion of system energy and, hence, are naturally related to control techniques for stabilization aiming to minimize the total energy. Since the minimum point of the Hamiltonian might not correspond to a desired regulation point, the control design needs to inject additional energy to ensure that the minimum of the total energy is at the desired equilibrium. For fully-actuated (port-)Hamiltonian systems, it is sufficient to shape the potential energy only using an energy-shaping and damping-injection (ES-DI) controller [31]. For underactuated systems, both the kinetic and potential energies needs to be shaped, e.g., via interconnection and damping assignment passivity-based control (IDA-PBC) [31, 54, 55, 56]. Wang and Goldsmith [57] extend the IDA-PBC controller from stabilization to trajectory tracking. Closely related to our controller design, Souza et. al. [58] apply this technique to design a controller for an underactuated quadrotor robot but use Euler angles as the orientation representation. Port-Hamiltonian structure and energy-based control design are also used to learn distributed control policies from state-control trajectories [59, 60, 61].
@@ IV-A Matrix Lie Groups
The adjoint $\mathsf{A}\mathsf{d}_{\boldsymbol{\mathfrak{q}}}:{\boldsymbol{\mathfrak{g}}}\rightarrow{\boldsymbol{\mathfrak{g}}}$ is defined as:
@@
The algebra adjoint $\mathsf{a}\mathsf{d}_{\boldsymbol{\mathfrak{q}}}:{\boldsymbol{\mathfrak{g}}}\rightarrow{\boldsymbol{\mathfrak{g}}}$ is the directional derivative of $\mathsf{A}\mathsf{d}_{\boldsymbol{\mathfrak{q}}}$ at ${\boldsymbol{\mathfrak{q}}}=\mathbf{e}$ in the direction of $\boldsymbol{\xi}\in{\boldsymbol{\mathfrak{g}}}$:
@@
The coadjoint $\mathsf{A}\mathsf{d}^{*}_{\boldsymbol{\mathfrak{q}}}:{\boldsymbol{\mathfrak{g}}}^{*}\rightarrow{\boldsymbol{\mathfrak{g}}}^{*}$ is defined as: $\langle\mathsf{A}\mathsf{d}^{*}_{\boldsymbol{\mathfrak{q}}}(\boldsymbol{\varphi}),\boldsymbol{\psi}\rangle=\langle\boldsymbol{\varphi},\mathsf{A}\mathsf{d}_{\boldsymbol{\mathfrak{q}}}(\boldsymbol{\psi})\rangle$. The algebra coadjoint $\mathsf{a}\mathsf{d}^{*}_{\boldsymbol{\xi}}:{\boldsymbol{\mathfrak{g}}}^{*}\rightarrow{\boldsymbol{\mathfrak{g}}}^{*}$ is the dual map of $\mathsf{a}\mathsf{d}_{\boldsymbol{\xi}}$, satisfying: $\langle\mathsf{a}\mathsf{d}^{*}_{\boldsymbol{\xi}}(\boldsymbol{\varphi}),\boldsymbol{\psi}\rangle=\langle\boldsymbol{\varphi},\mathsf{a}\mathsf{d}_{\boldsymbol{\xi}}(\boldsymbol{\psi})\rangle$.
@@ IV-C Hamiltonian Dynamics on Matrix Lie Groups
In this section, we describe Hamilton’s equations of motion on a matrix Lie group [62, 29]. Our neural network architecture design in Sec. V is based on Lie group Hamiltonian dynamics, which enables guarantees for both kinematic constraint satisfaction and energy conservation.
@@
$\dot{{\boldsymbol{\mathfrak{p}}}}$ $=\mathsf{a}\mathsf{d}^{*}_{\boldsymbol{\xi}}({\boldsymbol{\mathfrak{p}}})-\mathsf{T}_{\mathbf{e}}^{*}\mathsf{L}_{\boldsymbol{\mathfrak{q}}}\left(\frac{\partial{\cal H}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})}{\partial{\boldsymbol{\mathfrak{q}}}}\right)+\mathbf{B}({\boldsymbol{\mathfrak{q}}})\mathbf{u},$ (17b)
where $\mathsf{T}_{\mathbf{e}}^{*}\mathsf{L}_{\boldsymbol{\mathfrak{q}}}$ is the dual map of $\mathsf{T}_{\mathbf{e}}\mathsf{L}_{\boldsymbol{\mathfrak{q}}}$ and satisfies
$\langle\boldsymbol{\xi},\mathsf{T}_{\mathbf{e}}^{*}\mathsf{L}_{\boldsymbol{\mathfrak{q}}}(\boldsymbol{\eta})\rangle=\langle\mathsf{T}_{\mathbf{e}}\mathsf{L}_{\boldsymbol{\mathfrak{q}}}(\boldsymbol{\xi}),\boldsymbol{\eta}\rangle,$ (18)
for any $\boldsymbol{\eta}\in{\boldsymbol{\mathfrak{g}}}^{*}$ and $\boldsymbol{\xi}\in{\boldsymbol{\mathfrak{g}}}$. By comparing Eq. (13) and Eq. (IV-C), we have:

### v2

@@ Abstract
Accurate models of robot dynamics are critical for safe and stable control and generalization to novel operational conditions. Hand-designed models, however, may be insufficiently accurate, even after careful parameter tuning. This motivates the use of machine learning techniques to approximate the robot dynamics over a training set of state-control trajectories. The dynamics of many robots are described in terms of their generalized coordinates on a matrix Lie group, e.g. on $SE(3)$ for ground, aerial, and underwater vehicles, and generalized velocity, and satisfy conservation of energy principles. This paper proposes a port-Hamiltonian formulation over a Lie group of the structure of a neural ordinary differential equation (ODE) network to approximate the robot dynamics. In contrast to a black-box ODE network, our formulation embeds energy conservation principle and Lie group’s constraints in the dynamics model and explicitly accounts for energy-dissipation effect such as friction and drag forces in the dynamics model. We develop energy shaping and damping injection control for the learned, potentially under-actuated Hamiltonian dynamics to enable a unified approach for stabilization and trajectory tracking with various robot platforms.
@@ I Introduction
Motion planning and optimal control algorithms depend on the availability of accurate system dynamics models. Models obtained from first principles and calibrated over a small set of parameters via system identification [1] are widely used for unmanned ground vehicles (UGVs), unmanned aerial vehicles (UAVs), and unmanned underwater vehicles (UUVs). Such models may over-simplify or even incorrectly describe the underlying structure of the dynamical system, leading to bias and modeling errors that cannot be corrected by adjusting a few parameters. Data-driven techniques [2, 3, 4, 5, 6] have emerged as a powerful approach to approximate system dynamics with an over-parameterized machine learning model, trained over a dataset of system state and control trajectories. Neural networks are expressive function approximation models, capable of identifying and generalizing interaction patterns from the training data. Training neural network models, however, typically requires large amounts of data and computation time, which may be impractical in mobile robotics applications. Recent works [7, 8, 9, 10, 11, 12, 13, 14] have considered a hybrid (gray-box) approach, where prior knowledge of the physics, governing the system dynamics, is used to assist the learning process. The dynamics of physical systems obey kinematic constraints and energy conservation laws. These laws are known to be universally true but a black-box machine learning model might struggle to infer them from the training data, causing poor generalization. Instead, prior knowledge may be encoded into the learning model, e.g., using a prior distribution [3], a graph-network forward kinematic model [15], or a network architecture reflecting the structure of Lagrangian [16] or Hamiltonian [10] mechanical systems. Moreover, many physical robot platforms are composed of rigid-body interconnections and their state evolution respects the structure of a Lie group [17], e.g., the position and orientation kinematics of a rigid body evolve on the $SE(3)$ Lie group [18]. Existing works [19, 20, 21] on Lie group neural ODEs networks for learning dynamics and normalizing flows focus on preservation of the Lie group structure during backpropagation using an adjoint method, either via a higher-dimensional space [19, 20] or local coordinates [21, 22].
@@
Lagrangian and Hamiltonian mechanics [23, 24] provide physical system descriptions that can be integrated into the structure of a neural network [10, 25, 11, 26, 27, 28]. Prior work, however, has only considered vector-valued states, when designing Lagrangian- or Hamiltonian-structured neural networks. This limits the applicability of these techniques as many common robot systems have states on a Lie group. For example, Hamiltonian equations of motion are available for orientation but existing formulations rely predominantly on $3$ dimensional vector parametrizations, such as Euler angles [29, 30], which suffer from singularities. Similar to [19, 20], our work embeds matrix Lie groups in a higher-dimensional space $\mathbb{R}^{n\times n}$, allowing us to train the model via the widely used neural ODE network on Euclidean space [31]. However, we focus on embedding the Hamiltonian formulation of robot dynamics in the model and deriving a control policy for trajectory tracking. Concurrently, Wotte et al. [22] offer an exciting approach to learn $SE(3)$ Hamiltonian dynamics from data using a neural ODE on Lie group, trained via local coordinates [21] with guarantees of satisfaction of Lie group constraints. The authors designed an adjoint method on the Lie algebra for neural ODE training, and a potential shaping controller for stabilization of a fully actuated rigid body.
Our preliminary work [32] designed a neural ODE network [31] to capture Hamiltonian dynamics on the $SE(3)$ manifold [33], and derived a trajectory tracking control policy for potentially under-actuated systems. Our model is shown to provide accurate long-term trajectory predictions, respecting $SE(3)$ constraints and conserve total energy with high precision. Inspired by [27, 34], we model kinetic energy and potential energy by separate neural networks, each governed by a set of Hamiltonian equations on $SE(3)$. However, our preliminary work [32] is developed specifically for the $SE(3)$ manifold and does not model dissipation elements that drain energy from the system, such as friction or drag forces in real robot systems. The Hamiltonian formulation of robot dynamics can be generalized to port-Hamiltonian formulation, where the dynamics are governed by energy exchange, i.e., the law of energy conservation, energy dissipation, e.g., from friction, and energy injection, e.g., from control inputs. In this paper, we generalize our dynamics learning and control method in [32] using a port-Hamiltonian neural ODE network to embed general matrix Lie group constraints and introduce an energy dissipation term, represented by another neural network, to model friction and air drag in physical systems. We compensate for energy dissipation in the trajectory tracking control design to provide accurate tracking performance. We verify our approach extensively with simulated robot systems, including a pendulum and a Crazyflie quadrotor, and with several real quadrotor platforms. In summary, this paper makes the following contributions.
  • We design a neural ODE model that respects port-Hamiltonian dynamics over a matrix Lie group to enable data-driven learning of robot dynamics.
@@
  • We demonstrate our dynamics learning and control approach extensively with simulated robot systems (a pendulum and a quadrotor) and several real quadrotor robots.
@@ II Related Work
Hamiltonian-based methods [10, 25, 11, 26, 27, 28, 14] use a Hamiltonian formulation [23, 24] of the system dynamics, instead, in terms of generalized coordinates $\mathbf{{\boldsymbol{\mathfrak{q}}}}$, generalized momenta $\mathbf{{\boldsymbol{\mathfrak{p}}}}$, and a Hamiltonian function, $\mathcal{H}(\mathbf{{\boldsymbol{\mathfrak{q}}}},\mathbf{{\boldsymbol{\mathfrak{p}}}})$, representing the total energy of the system. Greydanus et al. [10] model the Hamiltonian as a neural network and update its parameters by minimizing the discrepancy between its symplectic gradients and the time derivatives of the states $(\mathbf{{\boldsymbol{\mathfrak{q}}}},\mathbf{{\boldsymbol{\mathfrak{p}}}})$. This approach, however, requires that the state time derivatives are available in the training data set. Chen et al. [11], Zhong et al. [27] relax this assumption by using differentiable leapfrog integrators [48] and differentiable ODE solvers [31], respectively. The need for time derivatives of the states is eliminated by back-propagating a loss function measuring state discrepancy through the ODE solvers via the adjoint method. Our work extends the approach in [27, 34] by formulating the Hamiltonian dynamics over a matrix Lie group, which enforces kinematic constraints in the neural ODE network used to learn the dynamics. Toth et al. [49] and Mason et al. [50] show that, instead of from state trajectories, the Hamiltonian function can be learned from high-dimensional image observations. Finzi et al. [26] show that using Cartesian coordinates with explicit constraints improves both the accuracy and data efficiency for the Lagrangian- and Hamiltonian-based approaches. In a closely related work, Zhong et al. [34] showed that dissipating elements, such as friction or air drag, can be incorporated in a Hamiltonian-based neural ODE network by reformulating the system dynamics in port-Hamiltonian form [51]. The continuous-time equations of motions in Lagrangian or Hamiltonian dynamics can also be discretized using variational integrators [52] to learn discrete-time Lagrangian and Hamiltonian systems [53, 54, 55, 56] and provide long-term prediction for control methods such as model predictive control [57]. This approach eliminates the need to use an ODE solver to roll out the dynamics but its prediction accuracy depends on the discretization time step. Meanwhile, our work encodes not only the Hamiltonian structure but also the Lie group constraints, satisfied by the states of rigid-body robot systems, such as UGVs, UAVs and UUVs, in a neural ODE network to learn robot dynamics from data.
While most existing dynamics learning approaches focus on Euclidean dynamics, many robot systems have states evolving on a matrix Lie group. Recent works on neural ODE networks [19, 20, 21, 22] on Lie groups are classified into either extrinsic [19, 20] or intrinsic methods [21, 22]. The extrinsic approach [19, 20] embeds a Lie group in a higher-dimensional space with Lie group constraints, enabling training with neural ODEs on Euclidean space. Our work belongs to the extrinsic approach by enforcing matrix Lie group constraints on its embedding space $\mathbb{R}^{n\times n}$, but focuses on incorporating the law of energy conservation, via a Hamiltonian formulation, in the dynamics model and providing control design for trajectory tracking. Meanwhile, the intrinsic approach [21, 22] develops adjoint methods for training using local coordinates on the Lie algebra, e.g. via local charts [21], and therefore, guarantees Lie groups constraints by design. Concurrently to our work, Wotte et al. [22] develop a neural ODE network on Lie group to learn Hamiltonian dynamics on the $SE(3)$ manifold by deriving an adjoint method on the Lie algebra, offering an exciting approach for learning structure-preserving dynamics model.
@@
While few dynamics learning papers consider control design based on the learned model, we develop a general trajectory tracking controller for Lie group Hamiltonian dynamics. The Hamiltonian formulation and its port-Hamiltonian generalization [51] are built around the notion of system energy and, hence, are naturally related to control techniques for stabilization aiming to minimize the total energy. Since the minimum point of the Hamiltonian might not correspond to a desired regulation point, the control design needs to inject additional energy to ensure that the minimum of the total energy is at the desired equilibrium. For fully-actuated port-Hamiltonian systems, it is sufficient to shape the potential energy only using an energy-shaping and damping-injection (ES-DI) controller [51]. For underactuated systems, both the kinetic and potential energies needs to be shaped, e.g., via interconnection and damping assignment passivity-based control (IDA-PBC) [51, 58, 59, 60]. Wang and Goldsmith [61] extend the IDA-PBC controller from stabilization to trajectory tracking. Closely related to our controller design, Souza et. al. [62] apply this technique to design a controller for an underactuated quadrotor robot but use Euler angles as the orientation representation. Port-Hamiltonian structure and energy-based control design are also used to learn distributed control policies from state-control trajectories [63, 64, 65].
@@ IV-A Matrix Lie Groups
The dot product definition above is used to define the dual maps in Def. 6 and 9, and loss functions in Sec. V-C and V-D.
@@
In this paper, we consider a matrix Lie group element ${\boldsymbol{\mathfrak{q}}}\in\mathsf{G}$ embedded in $\mathbb{R}^{n\times n}$, instead of its $n$-dimensional representations, e.g., $\log_{\mathsf{G}}^{\vee}{\boldsymbol{\mathfrak{q}}}$, due to potential issues of discontinuity [67] and singularity [68, 69]. For example, Zhou. et al. [67] show that higher-dimensional representations of rotations, e.g., $(n^{2}-n)$ dimensions for $SO(n)$, are more suitable for learning using neural networks because they ensure continuity.
@@
[…]
$\langle\mathsf{T}_{\mathbf{e}}^{*}\mathsf{L}_{\boldsymbol{\mathfrak{q}}}(\boldsymbol{\eta}),\boldsymbol{\xi}\rangle=\langle\boldsymbol{\eta},\mathsf{T}_{\mathbf{e}}\mathsf{L}_{\boldsymbol{\mathfrak{q}}}(\boldsymbol{\xi})\rangle,$ (6)
for any $\boldsymbol{\eta}\in{\boldsymbol{\mathfrak{g}}}^{*}$ and $\boldsymbol{\xi}\in{\boldsymbol{\mathfrak{g}}}$.
@@
For ${\boldsymbol{\mathfrak{q}}}\in\mathsf{G}$, the adjoint $\mathsf{A}\mathsf{d}_{\boldsymbol{\mathfrak{q}}}:{\boldsymbol{\mathfrak{g}}}\rightarrow{\boldsymbol{\mathfrak{g}}}$ is defined as:
@@
The algebra adjoint $\mathsf{a}\mathsf{d}_{\boldsymbol{\xi}}:{\boldsymbol{\mathfrak{g}}}\rightarrow{\boldsymbol{\mathfrak{g}}}$ is the directional derivative of $\mathsf{A}\mathsf{d}_{\boldsymbol{\mathfrak{q}}}$ at ${\boldsymbol{\mathfrak{q}}}=\mathbf{e}$ in the direction of $\boldsymbol{\xi}\in{\boldsymbol{\mathfrak{g}}}$:
@@
[…]
@@ IV-C Hamiltonian Dynamics on Matrix Lie Groups
In this section, we describe Hamilton’s equations of motion on a matrix Lie group [66, 33]. Our neural network architecture design in Sec. V is based on Lie group Hamiltonian dynamics, which encode both kinematic constraints and energy conservation.
@@
$\dot{{\boldsymbol{\mathfrak{p}}}}$ $=\mathsf{a}\mathsf{d}^{*}_{\boldsymbol{\xi}}({\boldsymbol{\mathfrak{p}}})-\mathsf{T}_{\mathbf{e}}^{*}\mathsf{L}_{\boldsymbol{\mathfrak{q}}}\left(\frac{\partial{\cal H}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})}{\partial{\boldsymbol{\mathfrak{q}}}}\right)+\mathbf{B}({\boldsymbol{\mathfrak{q}}})\mathbf{u}.$ (18b)
[…]
By comparing (14) and (18), we have: